### v1

# CRADMap: Applied Distributed Volumetric Mapping with 5G-Connected Multi-Robots and 4D Radar Sensing

Abstract
Sparse and feature SLAM methods provide robust camera pose estimation. However, they often fail to capture the level of detail required for inspection and scene awareness tasks. Conversely, dense SLAM approaches generate richer scene reconstructions but impose a prohibitive computational load to create 3D maps. We present a novel distributed volumetric mapping framework designated as CRADMap that addresses these issues by extending the state-of-the-art (SOTA) ORBSLAM3 [1] system with the COVINS [2] on the backend for global optimization. Our pipeline for volumetric reconstruction fuses dense keyframes at a centralized server via 5G connectivity, aggregating geometry, and occupancy information from multiple autonomous mobile robots (AMRs) without overtaxing onboard resources. This enables each AMR to independently perform mapping while the backend constructs high-fidelity 3D maps in real time. To overcome the limitation of standard visual nodes we automate a 4D mmWave radar, standalone from CRADMap, to test its capabilities for making extra maps of the hidden metallic object(s) in a cluttered environment. Experimental results Section-IV confirm that our framework yields globally consistent volumetric reconstructions and seamlessly supports applied distributed mapping in complex indoor environments. -Github Repository: http://surl.li/xgfpxd
I INTRODUCTION
The rapid development of robot autonomy has opened up new opportunities to automate routine tasks and improve safety in hazardous environments. With reliable simultaneous localization and mapping (SLAM) systems robust in pose estimation and map generation, there is a growing gap in multi-robot distributed SLAM to build detailed 3D maps using low-cost autonomous mobile robots (AMRs), with limited processing capabilities, for real-time and efficient mapping. Such systems are essential for applications such as navigation, path planning, and inspection. Mapping large complex indoor environments presents several unresolved problems. First, coordinating data from multiple robots to generate a dense, volumetric map requires significant computational power. Second, handling overlapping sensor data from multiple AMRs while maintaining distributed maps accuracy remains a critical issue. Third, although high-speed data transmission from WiFi offers great potential, managing high bandwidth with low latency for real-time performance is not trivial.
Industries such as manufacturing, warehouse management, and asset or infrastructure inspection, which rely on autonomous robotics in damage monitoring, digital inspections, and material handling, require the deployment of reliable solutions to enable the efficient completion of the aforementioned tasks. These environments are large and often cluttered with obstacles like machinery, walls, and structural elements, which can obstruct the view of visual sensors. This limitation affects tasks that require precise navigation and detailed mapping with a team of AMRs. As industrial applications become more demanding, there is a growing need for systems that can effectively manage these complexities and deliver reliable real-time mapping.
In response to these challenges we develop CRADMap, a novel scalable distributed volumetric mapping architecture that utilize ORBSLAM3 on the frontend [1] with the COVINS [2] on the backend, an overview shown in Figure. 1. In our approach, we enhance keyframes so that they not only store the camera pose (i.e., its position and orientation) and visual features but also include a complete 3D dense point cloud. This enriched data is transmitted over 5G to achieve high bandwidth with low latency to the backend server, which fuses keyframes from view, integrates dense point clouds, and reconstructs the scene to generate globally consistent, high-fidelity 3D maps in real-time. Additionally, since visual nodes miss objects hidden by obstacles, we automated an independent 4D mmWave radar that creates its point map to detect occluded or hidden metallic objects.
The main contributions of this work are as follows:
Mapping framework (CRADMap) that enhances Visual-SLAM with enriched keyframes containing dense point clouds to generate volumetric maps in real-time.
A multi-robot distributed architecture that leverages a high-bandwidth i.e. 5G network to offload intensive processing to a centralized COVINS backend, ensuring real-time, globally optimized 3D reconstructions.
Integration of an independent 4D mmWave radar that creates a standalone point cloud map, supplementing visual data to improve the detection of hidden structures in complex indoor settings.
Evaluation of live experiments from multiple AMRs to demonstrate mapping improvements 5G efficiency.
II RELATED WORK
II-A Volumetric mapping
Early RGB-D SLAM efforts by Henry et al. [4] and Endres et al. [5] achieved dense reconstructions yet often relied on offline optimization and high-end GPUs, making them impractical for real-time volumetric mapping. KinectFusion [6] introduced TSDF-based mapping for smaller indoor spaces, leveraging powerful hardware to sustain real-time performance. Later works such as voxel hashing by Nießner et al. [7] scaled KinectFusion to larger scenes, while Oleynikova et al. [8] (Voxblox) and Millane et al. [9] (Voxgraph) refined incremental fusion and drift correction mainly in single-AMR or offline settings. Additionally, an octree-based approach by Hou et al. [10] processed RGB-D video for real-time 3D indoor mapping, but it remained limited to a single-AMR pipeline without addressing the complexity of multi-robot collaboration. As a result, large-scale 3D volumetric maps on low-cost AMRs remain a key challenge, especially under live conditions.
II-B Multi-Robots SLAM Techniques and 5G Communication
Modern distributed and decentralized SLAM solutions such as Kimera-Multi [11] and Swarm-SLAM [12] extend multi-robot mappings frameworks advance multi-agent coordination but demand significant GPU resources on limited bandwidth networks. Meanwhile, CCM-SLAM [13] and COVINS [2] reduce global drift by merging keyframes on a centralized backend server, primarily handling sparse data. Our approach extends ORBSLAM3 [1] with COVINS [2], showing that centralized global optimization enhances keyframe accuracy for multi-robot in distributed architecture instead of map merging. To handle dense volumetric data from multiple robots in real-time, we utilize 5G communication, as highlighted by recent work [14] demonstrating how high throughput and low latency enable offloading of large sensor streams. By transmitting dense keyframes over 5G, our system alleviates each AMR computing load, allowing real-time distributed volumetric reconstruction.
II-C 4D mmWave Sensing Technology
Although radar-based SLAM [15], [16] can improve navigation in harsh visibility i.e. fog, snow, and low lighting, it introduces complex sensor fusion, and potential noise not present in indoor environments. Motivated from Doojin Lee et al [17] and Wallabot [18] illustrate the radar unique capability for detecting metallic or hidden objects behind walls. In our framework, radar operates separately from the CRADMap pipeline, focusing on detecting obscured objects that are not visible by camera or lidar. This modular design maintains a streamlined volumetric map referred to here as CRADMap and in addition, delivers a standalone radar point cloud map for detecting metal items that are obscured behind furniture and walls.
| Band | Name | Mode | FRaster (kHz) | Low (MHz) | Middle (MHz) | High (MHz) | DL/UL Bandwidth (MHz) |
| 78 | TD 3500 | TDD | 15, 30 | 3300 | 3550 | 3800 | 500 |
III METHODS
Our proposed system, CRADMap, is a distributed multi-robot framework for applied volumetric mapping that uses ORBSLAM3 [1] on front-ends, a centralized COVINS [2] on back-end for distributed architecture, and a dedicated Altos 4D mmWave radar [20] to generate additional point cloud map of obscured metallic object(s). Fig. 2 illustrates the detailed process blocks involved. The system leverages high-bandwidth 5G connectivity (24 ms end-to-end latency, up to 110 MBps) to enable four AMRs (autonomous mobile robots) as shown in Fig. 1. Each integrated with quectel 5G RM520N [21] to stream data in real-time with OAK-D pro [22] 3D RBG-D camera, with inherent scalability to AMRs. In the following, we describe our research pipeline supported by algorithms to illustrate the flow of implementation.
III-A Front-End: ORBSLAM3 with Dense Keyframe Generation
Each AMR processes an RGB-D data stream at time using ORBSLAM3 to extract features and estimate camera poses. Our method enriches standard ORBSLAM3 keyframes by generating dense point clouds, which facilitate robust volumetric mapping.
i. Feature Extraction and Pose Estimation
a) ORB Feature Extraction
For each frame , ORBSLAM3 extracts 2D keypoints (with indexing keyframes and indexing keypoints).
b) Depth Association
For each detected keypoint at pixel , the depth is retrieved from the depth image . The corresponding 3D point is computed via:
where is the back-projection operator using camera intrinsics ().
c) Pose Estimation
The camera pose is obtained by minimizing the reprojection error:
Here, projects a 3D point onto the image plane, ensuring alignment between observed key points and reprojected 3D points. SE(3) is a special Euclidean group in three dimensions.
ii. Dense Keyframe Generation
d) Dense Point Cloud Construction
When keyframe criteria (e.g., significant camera motion or scene change) are met, a dense point cloud is generated from the entire image:
where is the set of all pixel coordinates in image .
e) Outlier Removal
To improve robustness, we filter out noisy points by comparing each point to the local neighborhood statistics:
where and are the mean and standard deviation of neighboring points, and is a threshold. The cleaned point cloud is denoted as .
f) Keyframe Packaging
The keyframe is then stored as:
which includes the estimated pose , features, RGB-D images, and the cleaned dense point cloud.
iii. Image Plugins and 5G Bandwidth
For real-time streaming we use lossless PNG compression for RGB (15 FPS) and ZSTD for depth (10 FPS) ROS2 camera topics using image transport plugins [23]. Compressed images are sent to the cloud backend server for pose optimization and distributed map handling in COVINS. This integration ensures the creation of accurate dense volumetric maps. Each AMR is assigned a unique IPv6 address via the 5G Wwan, enabling real-time streaming and SLAM processing. We operate in the NR FR1 channel details presented in Table I which supports high data rates required for RGB-D cameras. The 5G network ensures fast IPv6-based discovery and low latency for experiments.
g) Bandwidth Calculation
each-AMR bandwidth ()
Total bandwidth for four AMRs:
III-B Back-End: COVINS for Optimizations
The COVINS back-end receives keyframes from all AMRs, stores them in a global Atlas, and performs loop closure and pose graph optimization to refine camera poses and reduce drift. COVINS improves Absolute Trajectory Error (ATE) and Root Mean Square Error (RMSE) through global consistency and loop closure, it maintain the individual maps from each AMR (RGB-D) in a distributed manner instead of collaborative map merging.
i. Keyframe Integration
a) Keyframes () are stored in the atlas as:
ii. Loop Closure and Global Pose Optimization
b) Loop Closure Constraint
For a new keyframe , candidate keyframes are identified via a Bag-of-Words approach. A loop closure constraint is computed as:
c) Global Pose Optimization
The global pose graph is refined by minimizing a cost function that combines odometry and loop closure constraints:
Here, represents the relative odometry error, and and are covariance matrices for odometry and loop closure constraints, respectively.
III-C Distributed Volumetric Mapping
In CRADMap, each AMR constructs its individual volumetric map from dense keyframes, transformed into the global coordinate frame using refined poses from COVINS. The distributed framework operates as follows: each AMR independently processes its RGB-D stream, generates dense keyframes from frontend (III-A), and builds a local map without map merging in a distributed manner (III-B). This pipeline ensures scalability by eliminating the need for computationally expensive map merging, and robustness against individual AMR failures. Each AMR generates a local map by transforming the dense point clouds from each keyframe into the global coordinate frame using the refined pose:
Then, the local map is formed as the union of all transformed points:
4D Radar-Metallic Object Detection
4D mmWave radar enable the detection of hidden metallic structures automatically triggers where camera fail. Weighting point cloud registration with SNR data filters outliers and filtering non-metallic returns, enhancing metallic detection for mapping hidden utilities as shown in section IV-D.
i. Radar Data Acquisition and SNR Filtering
At time , the radar produces a point cloud . Points are filtered based on SNR:
ii. Global Transformation and Noise Removal
Filtered radar points are transformed to the global frame:
A statistical outlier removal filter is then applied to :
IV Experiment
IV-A Evaluation Metrics and Dataset
We evaluate our system on the indoor settings (1) live experiments from the University of Waterloo (UW) E7 building and (2) TUM RGB-D benchmark [24]. In UW-E7 we captured 4 scenes as shown in Fig. 3 using four AMRs. All experiments are conducted on hardware Intel i7‑1355U, 32GB RAM, Graphics (RPL-U), and software Ubuntu 22.04, ROS2 Humble, with ORBSLAM3 [1] as our baseline. Volumetric mapping quality, CRADMap, and network performance are assessed on the UW-E7 dataset in Open3D [25] on campus Wifi eduroam [26] and 5G network provided by our industry partner, Rogers Inc [27]. 4D mmWave radar detection are evaluate by doing two complex experiments in UW-E7 shown in Fig. 4 and Fig. 5 where partial and complete block view. Lastly, pose estimation metrics performed on a public dataset.
IV-B CRADMap Evaluation
We generated four distributed dense volumetric maps on live UW-E7 experiments (Robohub Lab, Classroom, Stairs, and Faculty Corridor) using our CRADMap approach. To evaluate mapping, the qualitative comparison is shown in Fig. 3 also we use Open3D for quantitative evaluation to compute coverage (the percentage of the environment reconstructed via a voxel grid over the explored area) and density (points per cubic meter). First, the baseline ORBSLAM3 and our CRADMap are generated to ensure both are in the same coordinate frame, making our relative metrics robust. For example, in the Robohub Lab, glass surfaces cause ORB‑SLAM3 to achieve only about 26.91% coverage, while our method reconstructs approximately 78.93% of the environment with significantly higher point density as shown in Table LABEL:table:2. Similarly, in the Classroom, Stairs, and Faculty Corridor, our method consistently yields higher coverage and denser reconstructions than the baseline. These relative improvements in coverage 78% to 85% and 4.6–5.5 more density provide compelling evidence of our system’s enhanced mapping capabilities.
| Coverage (%) | Density (pts/m3) | ||||
| Scene | Baseline | Ours | Baseline | Ours | |
| Robohub Lab | 26.91 | 78.93 | 161 | 950 | |
| Classroom | 37.26 | 82.82 | 147 | 896 | |
| Stairs | 26.69 | 85.17 | 123 | 782 | |
| Faculty Corridor | 24.54 | 79.34 | 182 | 1023 | |
IV-C Network Performance Evaluation
Evaluation perform on live UW-E7 dataset in a multi-robot setup, where each AMR generates visual data at 14 MB/s (112 Mbps) from its RGB (15 Hz) and depth (10Hz) streams. The WiFi network at the UW employs IEEE 802.1X for authentication and both dual-band 2.4 GHz/5 GHz frequency bands. For the evaluation, the 5 GHz and WiFi band are utilized, a single AMR achieves an effective maximum upload speed of 90 Mbps, resulting in an update frequency of approximately 8.9 Hz; however, with 4 AMRs the total load reaches 56MB/s (448 Mbps), and the effective per AMR bandwidth drops to about 22.5 Mbps, which get reduced to around 3.5 Hz using 4 AMRs together. In contrast, 5G (3.5 GHz band, 110 to 120 Mbps upload) maintains a stable latency of 24 ms; with 1 AMR, the update frequency is about 10.4 Hz, and with 4 AMRs, it drops to 7.1 Hz due to shared bandwidth as shown in Table III. These calculations, based on the formula Map Update Frequency = (Effective Upload Speed per AMR)/(Data per Update), clearly show that 5G’s higher available bandwidth and stable latency yield significantly better and more predictable performance compared to WiFi.
| Data Rate (MB/s) | Performance | ||||
| Network | AMRs | Total | Latency (ms) | Map (Hz) | |
| WiFi | 1 | 14 | 8–33 (variable) | 8.9 | |
| 4 | 56 | 11–34 (variable) | 3.5 | ||
| 5G | 1 | 14 | 24 ± 2 | 10.4 | |
| 4 | 56 | 24 ± 4 | 7.1 | ||
IV-D Radar-Enhanced Occluded Object Detection
Behind-the-wall object detection enhance perception beyond visual occlusions [17]. To leverage this capability, we integrate camera-driven situational awareness to dynamically activate radar when a wall or occlusion region is detected using our Algorithm 3, ensuring efficient and targeted sensing. We evaluate 4D mmWave radar capabilities using a 4‑chip cascaded imaging radar operating at 77 GHz, which produces high-density point clouds with an angular resolution of 1.4 degrees in both azimuth and elevation. In the (1) scenario, an AMR maps a cluttered indoor aisle at UW-E7, as shown in Fig. 4, where the presence of furniture caused the CRADMap in Fig. 4(c) and baseline map in Fig. 4(d) to miss detecting a horizontal metal ventilation pipe. Missing such structural components can be dangerous, particularly for AMRs navigating environments, as undetected obstacles may cause collisions. The radar achieved a 92% ±3 detection rate for the pipe despite some residual noise. Another (2) scenario from Fig. 5(a,b), 3 cm thick, solid opaque wall made of plastic boards are used to completely block the view. Visibility beyond barriers is critical in scenarios such as search-and-rescue operations, and construction site inspections. Ground truth are establish via manual annotation placing marker points on the metal studs behind the wall and vent pipe as well, and after filtering out noise and non-metallic reflections, the radar system accurately detected all three metal studs, achieving a 84% ±5 detection rate. These results demonstrate that radar integration significantly enhances obscured object detection in environments where traditional visual nodes i.e. camera and lidar fail.
IV-E Pose Estimation Evaluation
Our primary goal is to perform volumetric mapping not improving the pose estimation but we evaluate on four TUM RGB-D sequences (fr1/plant, fr1/teddy, fr2/coke, fr2/dishes) in Table IV to validate consistency. We computed ATE and RMSE by averaging results over five runs using the available ground truth. Our method slightly improves ATE and RMSE compared to baseline because the COVINS backend enables global pose-graph optimization, aggregating loop closures to correct drift across the entire trajectory. Dense depth priors from volumetric fusion strengthen bundle adjustment in texture-less regions (e.g., walls, glass), where ORBSLAM3 sparse features struggles, leading to reduced feature matching in low-texture and reflective scenes. However, global pose-graph optimization and loop closures in the COVINS backend help mitigate drift in such scenarios.
| Sequence | Metric | ORBSLAM3 (m) ± | Ours (m) ± |
| fr1/plant | ATE | 0.1178 ± 0.0049 | 0.1043 ± 0.0045 |
| RMSE | 0.1323 ± 0.0052 | 0.1187 ± 0.0050 | |
| fr1/teddy | ATE | 0.1432 ± 0.0051 | 0.1275 ± 0.0048 |
| RMSE | 0.1654 ± 0.0060 | 0.1482 ± 0.0057 | |
| fr1/coke | ATE | 0.1012 ± 0.0043 | 0.0905 ± 0.0039 |
| RMSE | 0.1135 ± 0.0048 | 0.1116 ± 0.0044 | |
| fr1/dishes | ATE | 0.1476 ± 0.0059 | 0.1398 ± 0.0054 |
| RMSE | 0.1627 ± 0.0061 | 0.1635 ± 0.0056 |
V CONCLUSIONS
This work introduced CRADMap, a distributed volumetric mapping framework that bridges sparse and dense SLAM approaches for distributed multi-robot systems and detection in obscured environment. By enhancing ORBSLAM3 [1] with dense keyframes and leveraging COVINS [2] for backend optimization via 5G connectivity, CRADMap achieves globally consistent 3D reconstructions while offloading computational demands. The use of 5G ensures low-latency and high-bandwidth communication, enabling real-time data transmission and 2x faster map update. Experiments demonstrate significant improvements, with 75% to 85% environmental coverage and 4.6x to 5.5x higher point density compared to baseline. The automation of standalone 4D mmWave radar enables the detection of occluded metallic structures, validated in cluttered scenarios achieving detection rates of 84% to 92% where camera alone fail. Future work will investigate the scalability of mmWave 5G in larger environments and expand testing to outdoor scenarios. The primary limitation in our current evaluation is that it relies on relative comparisons with a baseline map, due to the absence of an external laser scanner.